### v4

# Geometry-Informed Neural Operator Transformer

Abstract
Machine-learning-based surrogate models offer significant computational efficiency and faster simulations compared to traditional numerical methods, especially for problems requiring repeated evaluations of partial differential equations. This work introduces the Geometry-Informed Neural Operator Transformer (GINOT), which integrates the transformer architecture with the neural operator framework to enable forward predictions on arbitrary geometries. GINOT employs a sampling and grouping strategy together with an attention mechanism to encode surface point clouds that are unordered, exhibit non-uniform point densities, and contain varying numbers of points for different geometries. The geometry information is seamlessly integrated with query points in the solution decoder through the attention mechanism. The performance of GINOT is validated on multiple challenging datasets, showcasing its high accuracy and strong generalization capabilities for complex and arbitrary 2D and 3D geometries.
keywords:
[…]
{graphicalabstract}[Uncaptioned image]
{highlights}
A Geometry-Informed Neural Operator Transformer (GINOT) is proposed for forward predictions on arbitrary geometries
GINOT encodes surface point clouds that are unordered, have non-uniform point density, and varying numbers of points.
GINOT effectively processes complex, arbitrary geometries and varying input conditions with high predictive accuracy
1 Introduction
Many natural phenomena in the physical world are governed by partial differential equations (PDEs), which encapsulate fundamental principles of physics. Numerical techniques such as finite difference, finite element, and finite volume methods have been developed to solve these equations computationally. While these methods enable high-resolution simulations to capture complex physical behaviors, they often demand significant computational resources, making them time-intensive, costly, and sometimes infeasible under current computational limitations.
To address these challenges, data-driven approaches, often referred to as surrogate models, have been developed. These methods utilize deep learning techniques to approximate PDE solutions, offering significant reductions in computational costs and faster simulations without compromising accuracy. Such advancements are particularly advantageous for applications requiring repeated PDE evaluations [1], including sensitivity analysis [2, 3], gradient-based optimization [4, 5], uncertainty quantification [6, 7], and inverse design enabled by generative AI [8, 9].
Neural operators have become prominent surrogate models, adept at mapping infinite-dimensional input functions to output functions at specified query points:
$G_{\theta}:\mathcal{F}\rightarrow\mathcal{G}(\textbf{x}),$ (1)
where ${G}_{\theta}$ is the neural operator with learnable parameter $\theta$, $\mathcal{F}$ is the space of input functions, and $\mathcal{G}$ is the space of output functions on query points x. Examples such as the Fourier Neural Operator (FNO) [10] and the Deep Operator Network (DeepONet) [11] have shown significant success in this area [12, 13, 14, 15, 16, 17, 18, 19, 20]. These models map inputs like initial and boundary conditions, material properties, and other functions to solution fields at query points. However, their applicability is often limited to simple geometries. For instance, FNO is tailored for fixed, regular meshes such as image grids or 3D voxels and requires a complete mesh graph as input due to its reliance on the Fast Fourier Transform kernel, making it unsuitable for arbitrary geometries and less scalable. While DeepONet can predict solution fields at arbitrary query points, its original formulation is restricted to single geometries with fixed query point locations and counts.
However, real-world problems frequently involve complex and arbitrary computational geometries. Recent advances, including graph neural operators (GNO) [21], Geo-FNO [22], GI-FNO [23], and Geom-DeepONet [24], have addressed challenges in handling diverse geometries. GNO can process irregular grids using graph representations but is limited to local operations under constrained computational budgets, resulting in lower accuracy. Both Geo-FNO and GI-FNO rely on projecting irregular meshes onto fixed regular grids and mapping them back. However, Geo-FNO requires a fixed number of query points during training, while GI-FNO constructs new graphs on regular meshes using local kernel integration layers through GNO, which need longer inference time. Geom-DeepONet, while overcoming the limitation of single geometries of DeepONet by removing or repeating nodes to ensure fixed data size, is restricted to parametric geometries (e.g., length, thickness, radius) and does not generalize to fully arbitrary shapes. Additionally, GI-FNO requires a signed distance function (SDF) as input to represent geometry, and Geom-DeepONet depends on both geometry parameters and the SDF. However, obtaining the SDF is not always easily, particularly for complex 3D geometries represented by surface meshes, as it incurs a computational cost of $O(N_{g}\times N_{m})$ for distance calculation and $O(N_{g}\times N_{m})$ for the minimum value, where $N_{g}$ is the number of grid points in the domain and $N_{m}$ is the number of surface mesh—both of which are large for intricate geometries.
Researchers have leveraged the attention mechanism of transformers [25] in neural operators to efficiently integrate information from input functions and query points [26, 27, 28, 29]. Neural Operator Transformers (NOT) are particularly advantageous for irregular meshes due to their attention mechanism, enabling the prediction of solution fields at arbitrary query points. Unlike DeepONet, which employs a simple dot product to combine information from input functions and query points, the attention mechanism in NOT directs each query point to focus on the most relevant information from the input functions. This capability allows NOT to generalize effectively and capture more complex relationships between input functions and query points. Building on this, the authors developed an SDF-based NOT that accommodates arbitrary geometries [9]. The SDF representation facilitates handling arbitrary geometries on fixed, regular grids, which are compatible with existing deep learning models. However, as noted earlier, obtaining the SDF is not always straightforward or computationally easily.
In this work, we introduce the Geometry-Informed Neural Operator Transformer (GINOT), which integrates the transformer architecture with the neural operator framework to enable forward predictions for arbitrary geometries without requiring SDF as geometry features. GINOT consists of two primary components: the geometry encoder and the solution decoder. Drawing inspiration from Shap-E [30] by OpenAI, we design the geometry encoder to process the boundary point cloud, representing the geometry, into the KEY and VALUE matrices for the attention mechanism. This choice facilitates the integration of information from query points in the solution decoder.
The geometry encoder utilizes sampling and grouping layers [31, 32] to extract local geometric features from the boundary point cloud. These local features are fused with global geometry information through cross-attention layers, enabling effective processing of unordered and non-uniformly distributed point clouds—unlike convolutional neural networks (CNNs), which operate on regular grids with uniform density. Because the number of points in the point cloud can vary between geometries, padding is applied to standardize input sizes within a batch. To ensure that padding points do not influence the geometry encoder’s output, a masking mechanism is employed: it prevents padding points from being selected in the sampling and grouping layers and excludes them from contributing to attention scores in the cross-attention layer.
In addition to varying geometries, many problems involve additional input parameters such as loading, material properties, and boundary conditions. To address such cases, we propose an extension strategy for GINOT by incorporating additional encoders to process these input functions.
This manuscript summarizes the key concepts, methodologies, and results of the GINOT model. We validate the effectiveness of GINOT on multiple challenging datasets and compare its performance with baseline methods. The results demonstrate that GINOT achieves high accuracy and generalizes effectively to arbitrary geometries.
2 GINOT
GINOT integrates the transformer architecture with the neural operator framework to enable forward predictions for arbitrary geometries. It consists of two primary components: the geometry encoder and the solution decoder, as illustrated in Fig. 1. Both components leverage the attention mechanism inherent in the transformer architecture [25], which facilitates the model’s ability to focus on relevant input information. The attention mechanism is defined as
$\text{Attention}(Q,K,V)=\text{softmax}\left(\frac{QK^{T}}{\sqrt{d_{e}}}\right)V,$ (2)
where $Q\in\mathbb{R}^{N_{q}\times d_{e}}$ represents the QUERY matrix, and $K,V\in\mathbb{R}^{N_{k}\times d_{e}}$ are the KEY and VALUE matrices, respectively. Here, $d_{e}$ denotes the embedding dimension, $N_{q}$ is the length of the query sequence, and $N_{k}$ is the length of the KEY and VALUE sequences. To ensure stable training, residual connections and layer normalization are applied within each attention block, as depicted in the bottom right of Fig. 1.
In the following sections, we detail the GINOT architecture, omitting the batch size dimension for simplicity. For instance, we describe the points cloud as having dimensions $N\times d$, where $N$ represents the number of points and $d$ denotes the dimensionality. In practice, the actual shape of the points cloud is $B\times N\times d$, with $B$ being the batch size.
Figure 1: Overview of the Geometry-Informed Neural Operator Transformer (GINOT) architecture. The boundary point cloud is initially processed through sampling and grouping layers to extract local geometric features. These local features are then fused with global geometric information via a cross-attention layer. This is followed by a series of self-attention layers and a final linear layer, producing the KEY and VALUE matrices for the cross-attention layer in the solution decoder. In the solution decoder, an multilayer perceptron (MLP) encodes the query points into the QUERY matrix for the cross-attention layer, which integrates the geometry information from the encoder. The output of the cross-attention layer is subsequently decoded into solution fields at the query points using another MLP.
2.1 Geometry encoder
Traditional convolutional architectures rely on highly regular input formats, such as image grids or 3D voxels, to efficiently perform kernel operations. While effective for structured data, these formats are unsuitable for capturing arbitrary geometries. In contrast, boundary points clouds offer a flexible representation for arbitrary geometries but lack the regularity required by convolutional architectures, presenting unique challenges for processing.
To address these challenges, we developed a geometry encoder that processes the boundary point cloud of a geometry into the KEY and VALUE matrices for the attention mechanism. This enables seamless integration of geometry information with query points in the solution decoder. In contrast to CNN inputs, boundary point cloud used to represent geometries have three key characteristics: (1) invariance to the order of points, (2) non-uniform point density across locations, and (3) a variable number of points across geometries. Drawing inspiration from Shap-E [30] by OpenAI, the geometry encoder, as shown in Fig. 1(a), is designed to handle these characteristics. It ensures invariance to point order and padding while maintaining robustness to variations in point density.
The point cloud is first processed through a sampling layer and a grouping layer for local geometry features, originally introduced in PointNet [31] and PointNet++ [32], which are well-suited to handle the aforementioned properties. The sampling layer employs the iterative farthest point sampling (FPS) method [32] to select $N_{s}$ points from the point cloud. This method begins by randomly selecting a point, then iteratively selects the farthest point from the already selected points until $N_{s}$ points are chosen. Compared to random sampling, the FPS method provides better coverage of the entire point set for a given number of centroids. The selected points are then passed to the grouping layer, which forms $N_{s}$ groups, each containing $N_{p}$ points.
For each group, the $N_{p}$ points are selected as those within a ball of radius $r$ (a hyperparameter) centered at the corresponding sampled point. If fewer than $N_{p}$ points exist within the ball, the group is padded with the nearest points to the center. Conversely, if more than $N_{p}$ points are present, the nearest $N_{p}$ points are chosen. Consequently, the output of the grouping layer has dimensions $N_{s}\times N_{p}\times d$, where $d$ represents the dimensionality of the point cloud. The sampling and grouping layers allow to extract local geometry features from nonuniform density point cloud [32]. These layers output the selected points along with their indices in the original point cloud, which are subsequently used in downstream layers.
The point cloud is also processed through a positional encoder layer utilizing the Nerf encoding method [33], followed by a linear layer. Using the indices from the sampling and grouping layers, the output is indexed to extract local geometry information, resulting in a tensor of shape $N_{s}\times N_{p}\times C$, where $C$ represents the output channel dimension of the positional encoder. This tensor is concatenated with the output of the grouping layer. The combined tensor is then passed through a series of 2D convolutional layers and a multi-layer perceptron (MLP), which aggregates the information within each group, reducing the dimension of $N_{p}$ to 1. As a result, the final output has a shape of $N_{s}\times C$.
The local information obtained by the sampling and grouping layers is utilized as the QUERY ($Q$) in the cross-attention blocks, while the KEY ($K$) and VALUE ($V$) are derived from the output of the positional encoder layer, which captures global geometry information. The attention mechanism allows the local information in the QUERY ($Q$) to focus on the relevant global geometry features. The QUERY ($Q$) can be invariant to the permutation of points order based on the sampling and grouping mechanism, whereas the KEY ($K$) and VALUE ($V$) depend on the order of points as they are directly derived from the positional encoder layer with the point cloud as input. However, the output of the cross-attention block remains invariant to the order of points because of:
$\text{Attention}(Q,K,V)_{id}=\text{softmax}\left(\frac{Q_{ij}K_{mj}}{\sqrt{d_{e}}}\right)V_{md},$ (3)
where the output is unaffected by permutations of the index $m$, which is a dummy index under Einstein summation notation.
The sampling and grouping layers, combined with the attention mechanism, allow that the geometry encoder handle the point cloud with varying points density, while remaining robust to the permutation of points order. Since the number of points in the point cloud varies across different geometries, padding is typically required to standardize input sizes within a batch for deep learning models. To prevent padding points from influencing the geometry encoder’s output, a masking mechanism is employed. This mechanism excludes padding points from being selected in the sampling and grouping layers. Additionally, padding points can interfere with attention scores in the cross-attention layer of the geometry encoder. To mitigate this, the corresponding entries in the attention score matrix $(Q\cdot K^{T})/\sqrt{d_{e}}$ are set to "$\infty$- ∞", ensuring that the attention layer’s output is unaffected by padding points. The flexibility and robustness of the proposed geometry encoder are illustrated in experiments Section 3.
2.2 Solution decoder
The solution decoder predicts the solution field at the query points using the geometry information from the geometry encoder. First, the query points are processed through a Nerf positional encoding, followed by an MLP, to generate the QUERY matrix with dimensions $N_{q}\times d_{e}$, where $N_{q}$ is the number of query points and $d_{e}$ is the embedding dimension. This QUERY matrix is then passed into cross-attention layers, which integrate the geometry information from the geometry encoder. The attention mechanism ensures that the query points focus on the most relevant geometry features. Finally, the output of the cross-attention layers is decoded into the solution field at the query points using another MLP. The architecture of the solution decoder is illustrated in Fig. 1(b).
Since the geometries are arbitrary, the number of query points varies across different geometries. To handle this variability, GINOT processes batches of geometries by padding the query points to match the maximum count within each batch, similar to techniques used for variable sequence lengths in natural language processing.
Figure 2: Extending GINOT for additional inputs. Encoders for extra inputs process these features, which are then concatenated with the geometry encoder’s output. The combined representation is passed through an MLP for aggregation and used as the KEY and VALUE in the solution decoder’s cross-attention blocks. This extension enables GINOT to incorporate both geometric and non-geometric inputs effectively.
2.3 Extending GINOT
Many problems involve not only varying geometries but also additional input parameters, such as loading conditions, material properties, or boundary conditions. To address such cases, GINOT can be extended by incorporating additional encoders to process these input functions. The outputs of these additional encoders are concatenated with the output of the geometry encoder and passed through an MLP for aggregation of all information from the geometries and the additional inputs. The aggregated output is then used as the KEY and VALUE in the cross-attention blocks of the solution decoder. The architecture of the extended GINOT is illustrated in Fig. 2. This extension is demonstrated in the bracket lug example in Section 3.4, where varying loading conditions are included as additional inputs.
3 Experiments
In this section, we present several numerical experiments to evaluate the performance and accuracy of GINOT on various challenging datasets. The primary evaluation metric used is the $L_{2}$ relative error defined as
$L_{2}=\frac{\|y^{\text{true}}-y^{\text{pred}}\|_{2}}{\|y^{\text{true}}\|_{2}},$ (4)
where $y^{\text{true}}$ and $y^{\text{pred}}$ represent the ground truth and predicted solutions, respectively. For some datasets, the $L_{2}$ relative error may not be an ideal metric. Therefore, we also employ the normalized root mean square error (NRMSE) as an additional metric defined as
$\text{NRMSE}=\frac{\sqrt{\frac{1}{N}\sum_{i=1}^{N}(y^{\text{true}}_{i}-y^{\text{pred}}_{i})^{2}}}{y^{\text{true}}_{\text{max}}-y^{\text{true}}_{\text{min}}}.$ (5)
To train the GINOT models, we use the mean square error (MSE) as the loss function. Since the number of query points is padded to match the maximum number of nodes in the batch, a masking mechanism is applied to exclude the padding points from the loss calculation:
$\text{MSE}=\frac{1}{1+\sum_{i=1}^{N}m_{i}}\sum_{i=1}^{N}m_{i}\left(y_{i}-\hat{y}_{i}\right)^{2},$ (6)
where $m_{i}$ is the mask value, set to 1 for non-padding points and 0 for padding points. The addition of 1 in the denominator ensures the denominator is never zero. Prior to training, both input and output data are normalized to have zero mean and unit variance.
This section provides an overview of the datasets used in the experiments, along with the training hyperparameters and computational efficiency of GINOT for these datasets. Additionally, we summarize GINOT’s performance on these datasets and compare it with baseline methods. Detailed experimental results are presented in the subsequent subsections. Further implementation details and results can be found in our GitHub repository.
3.1 Overview
To evaluate the performance and accuracy of GINOT, six datasets are utilized: elasticity [22], Poisson equation with both structured and unstructured meshes, bracket lugs [24], micro-periodic unit cell (PUC) [34] in [9], and Jet Engine Bracket (JEB) [35]. Table 1 provides a detailed summary of these datasets, including their size, the number of query points, the number of points in the boundary point cloud, whether the dataset is parametric or fully arbitrary, and the type of solution considered. The first three datasets have a constant number of points in the boundary point cloud, while the last three datasets feature a varying number of points in the boundary point cloud. All datasets feature non-uniform point densities, as illustrated in the visualization results in the following subsections.
Table 1: Dataset overview.
DataSet Elasticity[22] Poisson (structure mesh) Poisson (unstructure mesh) Bracket lug [24] Micro- PUC [34] JEB [35]
Data size 2000 6000 6000 3000 73879 2138
Number of query point 972 2737 [1566,2838] [5054,15210] [1136,6363] [18085,55009]
Number points in boundary PC 45 144 144 [2184,4284] [180,484] [2253,10388]
Parameteric True True True True False False
Solution Stress Mises stress Mises stress displacements Mises stress
Table 2 summarizes the training hyperparameters of the GINOT models. For the micro PUC dataset, two models are trained: one for predicting the solution at the final strain step and another for predicting the solution over historical strain steps.
Table 2: Training hyperparameters of the GINOT models. All models are implemented with the PyTorch scheduler ReduceLROnPlateau.
Elasticity Poisson (structure mesh) Poisson (unstructure mesh) Bracket lug Micro PUC (Final strain step) Micro PUC (historical strain steps) JEB
Batch size 20 32 32 32 64 64 32
Optimizer Adam Adam Adam Adam Adam Adam Adam
Initial learning rate 0.001 0.001 0.001 0.001 0.001 0.001 0.001
scheduler Patience 40 40 40 20 10 10 100
scheduler Factor 0.7 0.7 0.7 0.7 0.7 0.7 0.7
Epochs 1000 500 500 500 400 500 500
Training dataset 1000 80% 80% 80% 80% 80% 90%
Testing dataset 200 20% 20% 20% 20% 20% 10%
$N_{s}$ 64 64 64 512 128 128 512
$N_{p}$ 18 18 18 64 8 8 64
Grouping $r$ 0.2 0.2 0.2 0.5 0.2 0.2 0.5
Att. heads (Decoder) 4 8 8 8 4 4 1
Attention heads (Encoder) 4 8 8 8 8 8 1
Cross att. layers (Encoder) 1 1 1 1 2 2 1
Self att. layers (Encoder) 2 3 3 2 2 2 2
Cross att. layers (Decoder) 6 4 4 3 4 4 4
Table 3 summarizes the computational efficiency of GINOT. Training tasks were conducted on single NVIDIA A100 GPU on the DELTA machine or single NVIDIA H100 GPU on the DeltaAI machine, both hosted at the National Center for Supercomputing Applications (NCSA) at the University of Illinois Urbana-Champaign (UIUC). Inference tasks were executed on single NVIDIA A100 GPU on the DELTA machine.
Table 3: GINOT computation efficiency.
Elasticity Poisson (structure mesh) Poisson (unstructure mesh) Bracket lug Micro PUC (Final strain step) Micro PUC (historical strain steps) JEB
Training GPU A100 A100 A100 H100 H100 H100 H100
Training [s/epoch] 1.7 12.1 11.9 33.2 128 136 31
Inference [s/solution] 3.94e-4 1.03e-3 9.2e-4 1.35e-2 2.18e-3 4.82e-3 4.0e-2
Table 4 summarizes GINOT’s performance across the datasets, presenting the $L_{2}$ relative error or NRMSE for both training and testing samples. These results are compared against baseline methods, with the best performance highlighted in bold. The findings demonstrate GINOT’s ability to tackle more complex problems while outperforming the baseline methods. Detailed experimental results are provided in the subsequent subsections.
Table 4: Performance comparison of GINOT with baseline methods. “–" indicates that the method cannot handle the dataset, while “NA" signifies that the method can handle the dataset but results are unavailable. The best performance is highlighted in bold.
Dataset, Metric Data Geo-FNO[22] Geom- DeepONet[24] GNO[21] GI-FNO[23] GINOT
Elasticity, $L_{2}$ Training 1.25% NA 11.3% 1.32% 0.936%
Testing 2.29% NA 12.3% 1.66% 1.33%
Poisson, $L_{2}$ (structure mesh) Training 0.43% NA 7.43% 0.42% 0.39%
Testing 0.48% NA 7.82% 0.45% 0.437%
Poisson, $L_{2}$ (unstructure mesh) Training NA 7.98% 0.49% 0.41%
Testing NA 8.23% 0.55% 0.53%
Bracket lug,$L_{2}$ Training 1.818% 10.5% 2.51% 0.378%
Testing 1.824% 11.3% 2.81% 0.45%
Micro PUC, $L_{2}$ (Final strain step) Training 16.5% 9.24% 7.93%
Testing 18.2% 9.95% 9.05%
Micro PUC, $L_{2}$ (historical strain steps) Training 18.9% 9.56% 8.42%
Testing 19.5% 10.01% 9.06%
JEB, $L_{2}$ [NRMSE] Training 36.4% 27.5% [1.12%] 15.6% [0.768%]
Testing 67.3% 31.4% [1.34%] 35.6% [1.74%]
3.2 Elasticity
Figure 3: Stress solutions of the testing sample for the elasticity problem. The first row shows the median case ($L_{2}=1.23\%$), while the second row represents the worst case ($L_{2}=2.93\%$) based on $L_{2}$ relative error. The first column depicts the input surface point cloud, the second column presents the ground truth from finite element analysis, the third column shows the prediction by GINOT, and the last column highlights the absolute error between the prediction and the ground truth.
For the elasticity example, we utilized the dataset generated by Li et al. [22] for training the Geo-FNO model. The dataset consists of unit cells of $[0,1]^{2}$ with arbitrary-shaped voids at the center, as illustrated in Fig. 3. The central void is defined by 45 points evenly distributed along the tangential direction, parameterized as
$x_{i}$ $=r_{i}\cos(\theta_{i}),$ (7)
$y_{i}$ $=r_{i}\sin(\theta_{i}),$
where $r_{i}$ is the radius of the $i$-th point, and $\theta_{i}=(2\pi i)/45$. The radii $r_{i}$ follow a Gaussian distribution, constrained within $0.2\leq r_{i}\leq 0.4$. The unit cell is fixed at the bottom and a tensile traction is applied at the top to simulate the stress field. The dataset uses a structured mesh with a fixed number of 972 nodes across all samples. A total of 1000 training samples were used to train the model for predicting stress solutions for various geometries.
Geo-FNO [22] employs 45 parametric void radii as geometry features, which limits its applicability to parametric geometries. In contrast, GINOT uses boundary points clouds to represent geometry, enabling it to handle arbitrary geometries. The stress solutions of the testing samples predicted by GINOT are shown in Fig. 3, highlighting the median and worst cases in terms of $L_{2}$ relative error. For the median case, the $L_{2}$ relative error is 1.23%, while the worst case has an $L_{2}$ relative error of 2.93%. The mean $L_{2}$ relative error across the testing samples is 1.34%, with a standard deviation of 0.46%.
3.3 Poisson equation
Figure 4: Testing sample solutions for the Poisson equation. The first row shows the median case ($L_{2}=0.41\%$) while the second row presents the worst case ($L_{2}=2.27\%$) for unstructured mesh data, based on $L_{2}$ relative error. The third row displays the median case ($L_{2}=0.38\%$) for structured mesh data. The first column illustrates the input surface point cloud, the second column shows the FEM ground truth, the third column provides GINOT predictions, and the last column highlights the absolute error between predictions and the ground truth. Structured and unstructured meshes are also depicted in the last column.
The Poisson equation considered in this study is given by
$-\nabla^{2}u$ $=f,\quad\text{in}\quad\Omega,$ (8)
$u$ $=g,\quad\text{on}\quad\partial\Omega,$
where $\Omega$ represents the domain, $u$ is the solution, $f$ is the source term, and $g$ is the Dirichlet boundary condition. For all samples, we set $f=1$ and $g=0$. Similar to the elasticity example, the domain $\Omega$ is parameterized using Eq. 7, with the radius $r_{i}\in[0.2,0.8]$ and $i=1,2,\ldots,144$. The geometries are illustrated in Fig. 4.
Two datasets were generated for this problem, each containing 6000 samples. The first dataset uses a structured mesh with a fixed 2737 nodes across all samples, while the second dataset employs an unstructured mesh, where the number of nodes varies between 1566 and 2838. The structured and unstructured meshes are illustrated in the last column of Fig. 4. For comparison with Geo-FNO, which does not support unstructured meshes with varying node counts, we focus on the structured mesh dataset. The Poisson equation was solved using FEM with FEniCSX. Of the total samples, 20% were reserved for testing, and the remaining 80% were used for training.
The comparison between GINOT predictions and the ground truth is presented in Fig. 4. For the unstructured mesh dataset, the median case has an $L_{2}$ relative error of 0.41%, while the worst case has an $L_{2}$ relative error of 2.27%. For the structured mesh dataset, only the median case is shown for brevity, with an $L_{2}$ relative error of 0.38%. The mean $L_{2}$ relative error for the testing samples of the structured mesh dataset is 0.45%, with a standard deviation of 0.17%, while for the unstructured mesh dataset, the mean $L_{2}$ relative error is 0.44%, with a standard deviation of 0.21%.
3.4 Bracket lug
We address the plasticity problem of bracket lugs using the dataset used in [24]. The bracket lugs, depicted in Fig. 5, are parameterized by three variables: length, thickness, and the radius of the circular hole. FEM simulations employ an elastic-plastic material model with linear isotropic hardening and quadratic hexahedral elements under small deformation assumptions. For training GINOT, the middle edge nodes of the quadratic hexahedral elements are removed. The surface mesh is extracted from the volume hexahedral elements, and its nodes are used as the boundary point cloud, which is then input to the geometry encoder of GINOT.
The dataset also includes variations in pressure loading across samples. To incorporate the loading parameter, we adopt the framework shown in Fig. 2. An MLP encodes the loading (shape of $B\times 1$) into a tensor with shape of $B\times d_{e}$, which is expanded to $B\times N_{s}\times d_{e}$ by repeating the tensor $N_{p}$ times, where $B$ is the batch size. This tensor is concatenated with the geometry encoder’s output and passed through an MLP, whose output serves as the KEY and VALUE for the attention mechanism. The dataset contains 3000 samples, with 20% reserved for testing.
The stress solutions for the testing samples predicted by GINOT are shown in Fig. 5, highlighting the median and worst cases based on $L_{2}$ relative error. The median case has an $L_{2}$ relative error of 0.41%, while the worst case reaches 3.78%. The mean $L_{2}$ relative error for the testing samples is 0.45%, with a standard deviation of 0.23%.
Figure 5: Stress solutions for the testing samples of bracket lugs. The first row shows the median case ($L_{2}=0.41\%$), while the second row represents the worst case ($L_{2}=3.78\%$) based on $L_{2}$ relative error. The first column illustrates the input surface point cloud, the second column presents the ground truth from finite element analysis, the third column shows the GINOT prediction, and the last column highlights the absolute error between the prediction and the ground truth.
3.5 Micro-periodic unit cell
Figure 6: Mises stress and displacement solutions at the final strain step for the testing sample of the micro-periodic unit cell. The first row shows the median case ($L_{2}=10.37\%$), while the second row represents the 95th percentile case ($L_{2}=19.64\%$) based on $L_{2}$ relative error. The first column illustrates the input surface point cloud, the second column displays the true stress on the true deformed shape, the third column shows the predicted stress on the predicted deformed shape, and the last column highlights the absolute error of stress on the true deformed shape.
Figure 7: Mises stress and displacement solutions across historical strain steps for the median testing sample ($L_{2}=9.08\%$) of the micro-periodic unit cell, based on $L_{2}$ relative error. The first row shows the true stress on the true deformed shape, the second row presents the predicted stress on the predicted deformed shape, and the last row highlights the absolute error of the stress on the true deformed shape.
Figure 8: Demonstrating the versatility of the geometry encoder. The first row illustrates the input surface point cloud, with colors indicating the point order. The second row presents the corresponding predicted Mises stress on the predicted deformed shape by GINOT. (a) Original surface point cloud (PC) with 220 points, (b) shuffled original PC with the first 80% retained, (c) original PC padded with 199 points at coordinates (-1000, -1000) (not shown), and (d) shuffled version of the PC in (c). The $L_{2}$ relative errors for each case are displayed at the bottom.
The previous examples focus on 2D or 3D geometries with relatively simple shapes parameterized by a few variables. In this example, we consider micro-periodic unit cells (PUC) generated by the authors for metamaterial architecture design. These geometries are more complex and fully arbitrary. The dataset is created by first generating a 2D periodic Gaussian random field on a square domain $[0,1~{}\text{mm}]^{2}$. The Marching Squares algorithm is then applied to extract the contour at a random threshold value, defining the boundary of the 2D periodic unit cell, with field values below the threshold treated as void.
Periodic boundary conditions and compression are applied to the unit cell for FEM simulations to obtain the mechanical response field. The simulations use an elasto-plastic material model with large deformation within the environment of Abaqus. To approximate quasi-static compression while improving convergence for large deformation and elasto-plastic behavior, an implicit dynamic solver is employed with a virtual mass density of $\rho=10^{-8}$. The compression strain is applied smoothly over 51 steps from $\varepsilon=0$ to $\varepsilon=20\%$. The simulation utilizes a combination of mixed first-order quadrilateral elements and second-order triangular elements with full integration. To augment the dataset, the unit cell is randomly shifted multiple times. The corresponding stress and displacement solution fields are also shifted accordingly, enabling data augmentation without additional FEM simulations. The boundary point cloud of the geometry is formed by combining the points on the contours from the Marching Squares algorithm and the points on the boundary of the square domain, spaced at intervals of $\frac{1}{32}$ mm. This boundary point cloud represents the geometry features and is fed into the geometry encoder.
The dataset comprises around 73,879 samples, with 20% reserved for testing. The GINOT model is trained to predict Mises stress and two displacement components, both at the final strain step and over historical strain steps.
The predicted solutions at the final strain step are shown in Fig. 6, comparing the ground truth with GINOT predictions. The first row highlights the median case with an $L_{2}$ relative error of 10.37%, while the second row shows the 95th percentile case with an $L_{2}$ relative error of 19.64%. The first column illustrates the input surface point cloud, the second column presents the true stress on the true deformed shape, the third column displays the predicted stress on the predicted deformed shape, and the last column shows the absolute error of stress on the true deformed shape. The mean $L_{2}$ relative error across the testing samples is 9.05%, with a standard deviation of 3.53%.
The stress and displacement solutions over historical strain steps are compared with the ground truth and illustrated in Fig. 7. For brevity, only the median case is presented here. Additional cases, such as the best and worst, can be found in our Github repository. The first row shows the true stress on the true deformed shape, the second row depicts the predicted stress on the predicted deformed shape, and the last row highlights the absolute error of stress on the true deformed shape. The overall $L_{2}$ relative error for historical steps is 9.08%. The mean $L_{2}$ relative error for the testing samples is 9.06%, with a standard deviation of 3.02%.
As detailed in Section 2.1, the geometry encoder in GINOT is designed to handle 1) nonuniform point cloud densities, 2) permutations in point order, and 3) padding within the point cloud. The first property is evident throughout the examples in this work. To demonstrate the latter two properties, we analyze the second-best test case in this example. Notably, GINOT maintains stable inference performance even when the point cloud is made sparser. In this analysis, the point cloud is shuffled, reduced to 80% of its original size, padded with 199 points at $(-1000,-1000)$, and shuffled again. These modified point clouds are input to the trained GINOT, with results shown in Fig. 8. The $L_{2}$ relative errors for the original, reduced and shuffled, padded, and shuffled padded point clouds are 4.24%, 4.65%, 4.24%, and 4.57%, respectively. The similar $L_{2}$ errors across these cases demonstrate the geometry encoder’s robustness to point order and padding. However, a substantial reduction in point cloud density can increase the $L_{2}$ error due to loss of geometric information, while shuffling and padding have negligible impact.
Figure 9: Workflow for the Jet Engine Bracket (JEB) dataset. (a) Boundary points clouds are extracted from the CAD models of each JEB sample using FreeCAD with PythonAPI and serve as input to the geometry encoder of GINOT. The 3D volume mesh nodes are utilized as query points and provided to the solution decoder of GINOT. (b) Distribution of the number of boundary points in the cloud and volume mesh nodes across the 2138 JEB samples.
Figure 10: Training and validation mean squared error (MSE) loss curves for GINOT on the JEB dataset.
3.6 Jet engine bracket
The final example demonstrating GINOT’s capability is the Jet Engine Bracket (JEB) dataset [35]. This dataset comprises 2,138 samples of jet engine brackets derived from the "GE Jet Engine Bracket Challenge" guidelines. These geometries are complex 3D structures with arbitrary shapes and sizes. FEM simulations provide solution fields (e.g., stress, displacement) under four loading scenarios: vertical, horizontal, diagonal, and torsional, all with fixed loading magnitudes. In this example, we focus on the Mises stress solution under vertical loading.
Figure 11: Mises stress solutions for the testing samples of the JEB dataset. The first row shows the best case ($\text{NRMSE}=0.74\%$, $L_{2}=16.9\%$), the second row represents the median case ($\text{NRMSE}=1.62\%$, $L_{2}=31.2\%$), and the third row depicts the worst case ($\text{NRMSE}=3.54\%$, $L_{2}=56.8\%$) based on NRMSE. The first column illustrates the input surface point cloud, the second column presents the ground truth from finite element analysis, the third column shows the GINOT prediction, and the last column highlights the absolute error between the prediction and the ground truth.
The dataset includes CAD models, volume meshes in VTK format, and solution fields on fine quadratic tetrahedral elements. Instead of extracting surface nodes from the volume mesh, we extract the boundary point cloud directly from the CAD model vertices for each JEB sample using FreeCAD with PythonAPI, as illustrated in Fig. 9(a). To reduce computational costs, the volume mesh is downgraded from 10-node tetrahedrons to 4-node tetrahedrons by removing the middle nodes of each element’s edge. The nodes of these linear tetrahedrons are then used as query points in the solution decoder of GINOT, as shown in Fig. 9(a). Fig. 9(b) illustrates the number of boundary points in the cloud and the volume mesh nodes for the 2,138 JEB samples. The boundary point cloud contains between 2,253 and 10,388 points, while the volume mesh nodes range from 18,085 to 55,009.
The dataset is divided into 90% for training and 10% for testing. Fig. 10 illustrates the training process of GINOT on the JEB dataset. The mean squared error (MSE) loss for both the training and validation datasets decreases steadily during training. However, the testing loss remains significantly higher than the training loss, suggesting overfitting. This behavior is expected given the complexity of the JEB dataset and its relatively small size.
After training, we evaluate GINOT’s performance on both the training and testing samples. For the training samples, the mean $L_{2}$ relative error is 15.6% with a standard deviation of 3.6%, while for the testing samples, the mean $L_{2}$ relative error is 35.6% with a standard deviation of 9.88%. We also assess performance of the JEB dataset using the normalized root mean square error (NRMSE) defined in Eq. 5. The mean NRMSE for the training samples is 0.768% with a standard deviation of 0.167%, and for the testing samples, the mean NRMSE is 1.73% with a standard deviation of 0.525%. Fig. 11 illustrates the Mises stress solutions for the testing samples predicted by GINOT, showcasing the best, median, and worst cases based on NRMSE. The best case achieves an NRMSE of 0.74% and $L_{2}=16.9\%$, while the worst case has an NRMSE of 3.54% and $L_{2}=56.8\%$.
4 Conclusions
In this work, we introduced the Geometry-Informed Neural Operator Transformer (GINOT), a novel framework for efficient and accurate solution prediction on arbitrary geometries represented by unordered, non-uniform, and variable-sized surface point clouds, without the need for additional geometric features such as signed distance functions (SDFs). By utilizing transformer-based attention mechanisms and a geometry-aware encoding strategy, GINOT effectively captures geometric representations from boundary point clouds and maps them to solution fields at arbitrary query points. Unlike existing neural operator methods constrained by structured meshes, fixed query points, or computationally intensive geometric representations, GINOT offers a flexible and scalable approach for solving complex geometry problems. To enhance its functionality, we have integrated additional encoders to process problem-specific inputs, such as material properties and boundary conditions.
Extensive experiments on diverse datasets have demonstrated GINOT’s excellent accuracy and strong generalization capabilities across various geometries. These results establish GINOT as a powerful surrogate model for computational physics involving complex domains, providing significant computational speedups while maintaining high predictive accuracy. Future work will focus on extending GINOT to tackle multi-physics problems, further enhancing its utility in scientific computing and engineering applications.
5 Data availability
The dataset is available on Zenodo [36] and the trained models are available at the GitHub repository https://github.com/QibangLiu/GINOT
6 Code availability
The codes for training and inference are available at the GitHub repository at https://github.com/QibangLiu/GINOT
7 Acknowledgements
The authors would like to thank the National Center for Supercomputing Applications (NCSA) at the University of Illinois, and particularly its Research Computing Directorate, Industry Program, and Center for Artificial Intelligence Innovation (CAII) for support and hardware resources. This research used both the DeltaAI advanced computing and data resource, which is supported by the National Science Foundation (award OAC 2320345) and the State of Illinois, and the Delta advanced computing and data resource which is supported by the National Science Foundation (award OAC 2005572) and the State of Illinois. Delta and DeltaAI are joint efforts of the University of Illinois Urbana-Champaign and its National Center for Supercomputing Applications.
8 Author contributions statement
Q. Liu conceptualized the study, developed the methodology, implemented the models, prepared the datasets, conducted the analysis, and drafted, reviewed, and edited the manuscript. W. Zhong contributed to running the baseline models (GNO and GI-FNO), manuscript review and editing. H. Meidani contributed to conceptualization, results discussion, and manuscript review and editing. D. Abueidda contributed to results discussion and manuscript review and editing. S. Koric prepared the bracket lug dataset, contributed to conceptualization, results discussion, and reviewed and edited the manuscript. P. Geubelle contributed to the supervision, methodology, results discussion, and manuscript review and editing.
9 Competing interests statement
The authors declare no competing financial or non-financial interests.